At what interest rate (compounded annually) will a sum of $6000 grow to $8000 in 3 years? (Round your answer to two decimal places.)

Answer by mananth(16946)

i= (fv/pv)^(1/n) -1

fv= 8000
pv= 6000
n= 3
i=( 8000 / 6000 )^ 1 / n - 1
i= 1.33 ^ 1 / 3 - 1
i= 0.07
i= 7 %